OpenStudy (stamp):

15x - 12y + 3 = 0, Graph

OpenStudy (stamp):

\[15x-12y+3=0\]\[15x+3=12y\]\[15x/12+3/12=12y/12\]\[5x/4+1/4=y\]

OpenStudy (anonymous):

good whats next